OpenStudy (anonymous):

What is 7.8 × 105 m/h in kilometers per day? A. 3.24 × 1010 km/day B. 3.24 × 108 km/day C. 1.872 × 106 km/day D. 1.872 × 104 km/day

OpenStudy (anonymous):

@PaulaLovesSchool13

OpenStudy (anonymous):

@I_Need_Help_With_Home

OpenStudy (anonymous):

First you know that 1 m = 1/1000 km and 1h = 1/24 day then you can apply these equation

OpenStudy (anonymous):

If you apply it all to the only one that is correct is 1.872 × 104 km/day

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

Your welcome